OpenStudy (anonymous):

Rational Functions Help: A box has a volume of x^3-16x cm^3. Another box has a volume of x^3-4x^2 cm^3. What is the ratio of the volume of the first box to that of the second? What is the sum 3x/x^2-9 + 3/x^2-6x+9? State any restrictions on the variable. What is the difference 4x/x^2-3x+2 - 5/6x-12 State any restrictions on the variable. @whpalmer4

OpenStudy (mathstudent55):

For the first part, since you need a ratio, write the two expressions as a fraction, with the first expression in the numerator, and the second expression in the denominator. Then factor the numerator and denominator. What do you get?

OpenStudy (whpalmer4):

For all of these, you might as well factor numerator and denominator, as that will ease your job of making a common denominator. To find restrictions on the variable, find all values of \(x\) that make the denominator equal 0. \(x\) cannot take on those values, as doing so would result in a division by 0, which has an undefined value.

OpenStudy (anonymous):

@mathstudent55 After simplifying the ratio and factoring it. I got x(x+4)/x^2. @whpalmer4 How would I begin factoring The sum and the Difference of problems though? :/

OpenStudy (anonymous):

@whpalmer4 Can you help me set up the second problem because I am so confused?

OpenStudy (whpalmer4):

What is the sum \[\frac{3x}{x^2-9} + \frac{3}{x^2-6x+9}\] As I suggested, you need to factor the denominators. Do you know how to factor the expressions \[x^2-9\]and\[x^2-6x+9\]

OpenStudy (anonymous):

(x+3) and (x-3) (x-3) (x-3) After that Idk what to do.

OpenStudy (whpalmer4):

Okay, that's good. Now write the fractions as factored:\[\frac{3x}{(x-3)(x+3)} + \frac{3}{(x-3)(x-3)}\] Now before you go any further, figure out which values of \(x\) will cause the denominator of either fraction to equal 0. Those values of \(x\) are the restricted values.

OpenStudy (whpalmer4):

Next, establish a common denominator for both fractions. Do you know how to do that?

OpenStudy (mathstudent55):

First Part. Here is the ratio of the two volumes written as a fraction. \(\dfrac{x^3-16x}{x^3-4x^2} \) Now we factor the numerator and denominator. The first step is to factor out a common factor. Both the numerator and the denominator have common factors. \(=\dfrac{x(x^2-16)}{x^2(x-4)} \) Now we factor \(x^2 - 16\) in the numerator. As you wrote above, you know this is the difference of two squares and factors into \(x + 4)(x - 4) \). \(=\dfrac{x(x + 4)(x - 4)}{x^2(x-4)} \) Now we divide the numerator and denominator by the common terms (which is usually referred to as canceling out common terms.) \(=\dfrac{\cancel{x}(x + 4)\cancel{(x - 4)}}{\cancel{x}x\cancel{(x-4)}} \) We are left with \(= \dfrac{x + 4}{x} \) Your answer was \(\dfrac{x(x + 4)}{x^2} \). Your answer is equivalent to the given problem and to the final answer, but you did not reduce the fraction completely because you could have divided the numerator and denominator by x. In other words, your answer is a correct step in trying to find the final answer.

OpenStudy (anonymous):

@mathstudent55 Okay thank you so much!! So my final answer would be x+4/x? Do I have to cube the answer to find the ratio of the volumes? Because it asks "What is the ratio of the volume of the first box to that of the second?" @whpalmer4 So the restrictions would be x cannot equal 3 or -3? And a common denominator I think would be (x+3)(x-3)^2. Right?

OpenStudy (mathstudent55):

No. Don't cube anything. That is the final answer for the first problem. Each expression you were given was already a volume. It just so happened that the expressions, once written as a ratio in a fraction, could be simplified. If you write your answer is a single line, then it should be written as (x + 4)/x (using parentheses around the x + 4.)

OpenStudy (whpalmer4):

Yes, \((x+3)(x-3)^2\) would be the logical common denominator. Convert both fractions to that denominator and combine, then simplify if possible.

OpenStudy (whpalmer4):

Also, those are the correct restrictions on \(x\).

OpenStudy (anonymous):

@mathstudent55 Oh okay thank you soo much! How do you award medals on here? you and @whpalmer4 were great help!! :)

OpenStudy (whpalmer4):

Click the blue "Best Response" button on a post that helped you. Only can do one per question, go ahead and give it to @mathstudent55, he typed more :-)

OpenStudy (mathstudent55):

Click on Best Response, but you can only give one medal per post. I gave you one, so give one to whpalmer4.

OpenStudy (anonymous):

Alright, thank you to both of you! I am new on here so I am still trying to figure this thing out. How do you convert both fractions and combine? @whpalmer4

OpenStudy (whpalmer4):

A favorite saying of mine: "it's amazing what can be accomplished when no one cares about who gets the credit" :-)

OpenStudy (whpalmer4):

Multiply by a fraction that equals 1, but is written as missing terms / missing terms: so if you have \[\frac{(x-2)}{(x+3)}\]and you want it over \((x+3)(x-3)\) you would multiply by this: \[\frac{(x-2)}{(x+3)}*\frac{(x-3)}{(x-3)} = \frac{(x-2)(x-3)}{(x+3)(x-3)} \] then expand the numerator and add with the other fraction(s)

OpenStudy (whpalmer4):

(just a made-up example, any similarly to the problem at hand is strictly coincidental!)

OpenStudy (whpalmer4):

similarity, not similarly...

OpenStudy (whpalmer4):

really, it's just the same as if you were operating on numbers alone...

OpenStudy (anonymous):

Alright since the LCD is (x+3)(x-3)^2? How would I do the square to get it for the denominators of both fractions? I am sorry, I'm just so confused

OpenStudy (anonymous):

Square the 3x? so 3x^2/(x+3)(x-3)^2?

OpenStudy (whpalmer4):

\[(x+3)(x-3)^2 = (x+3)(x-3)(x-3)\] Your left hand fraction is \[\frac{3x}{(x-3)(x+3)} \]so we multiply it by \[\frac{(x-3)}{(x-3)}\] Your right hand fraction is \[\frac{3}{(x-3)(x-3)}\]so we multiply it by \[\frac{(x+3)}{(x+3)}\] when you do that and expand the numerators, you should be able to add them together and simplify (might have to factor the numerator again to do so)

OpenStudy (whpalmer4):

in both cases, we multiply by a fraction with the value 1, so we aren't doing anything except changing to a more convenient form.

OpenStudy (whpalmer4):

It's also possible that after you combine them that you won't be able to appreciably simplify, but you won't know until you do the work.

OpenStudy (whpalmer4):

However, as I mentioned, it is important to do the check for restricted values first, before any cancellation of common terms from numerator and denominator, because the restrictions still apply, even if the terms that contributed them appear to go away.

OpenStudy (anonymous):

Okay, So i got 3x^2-9/(x-3)(x+3)(x-3) and 3x^2 +9/(x-3)(x+3)(x-3). is that right?

OpenStudy (whpalmer4):

For example\[\frac{x(x+2)}{(x+2)(x-2)} =\frac{x}{x-2} \]but it still has the restriction \[x \ne -2\] as well as \[x \ne 2\]

OpenStudy (whpalmer4):

Uh, no you need to recheck your work. Also, if you are going to write a fraction on one line like that, you must put parentheses around the numerator and denominator. What you've written is actually \[3x^2 - \frac{9}{(x-3)(x+3)(x-3)} + 3x^2 + \frac{9}{(x-3)(x+3)(x-3)}\]

OpenStudy (whpalmer4):

(3x^2-9)/(x-3)(x+3)(x-3) would be the way to write the left hand fraction on one line

OpenStudy (whpalmer4):

actually, (3x^2-9)/((x-3)(x+3)(x-3)) would be safer.

OpenStudy (whpalmer4):

just doing the numerators: \[3x*\frac{(x-3)}{(x-3)} = \] \[3*\frac{(x+3)}{(x+3)} = \] See your mistakes?

OpenStudy (anonymous):

oh so the right hand fraction is 3x+9/(x-3)(x-3)(x+3) I hope that was right :/

OpenStudy (whpalmer4):

Still didn't put parentheses around your numerator! Also, you need to go recheck your work on the left fraction.

OpenStudy (anonymous):

Okay so it would be (3x+9)/...... I guess I am just still confused where did I mess up at? You said the left fraction was right, right? Just the right fraction was wrong?

OpenStudy (whpalmer4):

No, I didn't say the left was correct.

OpenStudy (anonymous):

Alright sorry about that, I'll re-do these problems again.

OpenStudy (whpalmer4):

Just check your work carefully. \[3x*\frac{(x-3)}{(x-3)} = \]

OpenStudy (anonymous):

(3x^2-9x)/(x-3). If you factor it you get 3x(x-3)/(x-3) and the two (x-3)'s cancel out giving you 3x?

OpenStudy (whpalmer4):

Okay, you don't want simplify here, you're just constructing the terms to add for the numerator. You did spot your mistake, I see, good :-) \[\frac{3x^2-9x}{(x-3)(x+3)(x-3)} + \frac{3x+9}{(x-3)(x+3)(x-3)} = \]

OpenStudy (anonymous):

Alright :) So do I simplify the numerators at this point and start canceling like terms?

OpenStudy (whpalmer4):

Combine the two fractions, then simplify, if possible.

OpenStudy (anonymous):

So it would be (3x^2-6x+9)/(x-3)(x+3)(x-3) after combining?

OpenStudy (whpalmer4):

Again, you really need to put parentheses around that denominator if you want it to mean the right thing. Yes, you have the right combined fraction. Unfortunately, the only factoring that can be done at this point is to factor out a 3 from the numerator, which doesn't really simplify the fraction at all.

OpenStudy (anonymous):

What do you mean? I thought I did put parentheses in the denominator? I put (x-3)(x+3)(x-3). So it would be 3(x^2-2x+3)/(x-3)(x+3)(x-3) as my final answer since it cant be further simplified?

OpenStudy (whpalmer4):

No, you need parentheses around all of those terms as well: \[3(x^2-2x+3)/((x-3)(x+3)(x-3))\]otherwise what you are writing is \[\frac{3(x^2-2x+3)}{(x-3)}(x+3)(x-3)\]

OpenStudy (anonymous):

Oh, okay! So 3(x^2-2x+3)/((x-3)(x+3)(x-3)) is the final answer since I cannot simplify anymore? Because i don't see anymore ways to simplify this answer.

OpenStudy (whpalmer4):

Yes. If you wanted to make it a little tidier, you could of course write it as \[\frac{3(x^2-2x+3)}{(x+3)(x-3)^2}\]or 3(x^2-2x+3)/((x+3)(x-3)^2)

OpenStudy (anonymous):

Alright, thank you so much. That is starting to make a little more sense. Don't kill me, but could you help me through the 3rd problem if I tried doing it?

OpenStudy (anonymous):

For the 3rd one I got (19x-5)/(6(x-2)(x-1)). Is that right? or do I need to re check my work?

OpenStudy (whpalmer4):

Assuming the 3rd problem is supposed to be \[\frac{4 x}{x^2-3 x+2}-\frac{5}{6 x-12}\]I do not get the same answer. Also, what are your restricted values?

OpenStudy (anonymous):

I got x does not equal 2 and 1.

OpenStudy (whpalmer4):

agree on the restricted values. try your fraction algebra again

OpenStudy (anonymous):

Okay it should say 4x/((x-2)(x-1)) - 5/(6(x-2)). The LCD would be 6(x-2)(x-1)?

OpenStudy (whpalmer4):

yes hint: we got the same results, but you had different signs in the numerator. I'll check back later tonight, but I'm going offline for 5 hours or so.

OpenStudy (anonymous):

Okay I will re-do that problem! Thanks again for your help! :)

OpenStudy (anonymous):

Would the answer be (19x+5)/(6(x-2)(x-1))? Sorry for responding so late, internet has been down

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Yes, you found your mistake (or at least didn't repeat it!)

OpenStudy (anonymous):

Okay so the final answer is (19x+5)/(6(x-2)(x-1)) with restrictions on 2 and 1?

OpenStudy (whpalmer4):

It is, or my name isn't Mortimer J. Schnicklegruber III :-)

OpenStudy (anonymous):

That's a pretty awesome name ;) Thank you so much! In fact, I have a surprise for you! Just follow along with my plan :)